Ruby is 14 years older than her little brother Harry. In six years she will be twice as old as Harry. How old is Ruby now?
Answer by Maths68(1474)

Ruby’s Present Age = R
Harry’s Present Age = H
Ruby is 14 years older than Harry. If we add 14 years to Harry’s age both Harry and Ruby will be of same age
R=H+14 …………(1)
In six years means we should add 6 to the Ruby and Harry ages.
R+6=H+6
In six years Ruby will be twice as old as Harry means
R+6=2(H+6) ………….(2)
Now solve (2)
R+6=2(H+6)
R+6=2H+12
R=2H+12-6
R=2H+6 …………..(3)
Put the value of R from equation (1) to equations (3)
H+14=2H+6
14-6=2H-H
8=H
H=8
Put the value of H in equation (1)
R=8+14
R=22
So Ruby is 22 and Harry is 8 years old.